OpenStudy (anonymous):

PLEASE HELP SIMPLYFIYING! 1/1+x/4 i know the answer when simplified is 4/x+4 but I don't know how to get that. thank youu

OpenStudy (anonymous):

\[\frac{ 1 }{ 1+\frac{ x}{ 4 } }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

first you have in the denominator \[1+\frac{ x }{ 4 }=\frac{ 4+x }{ x }\]

OpenStudy (anonymous):

sorry \[1+\frac{ x }{ 4 }=\frac{ x+4 }{ 4 }\]

OpenStudy (anonymous):

so the final ans is \[\frac{ 4 }{ 4+x }\]

OpenStudy (anonymous):

oh so that's all i had to do?

OpenStudy (anonymous):

yea